OpenStudy (anonymous):

Consider the solid obtained by rotating the region bounded by the given curves about the line x=2 x=2y^2,x=2 Find the volume V of this solid I'm having a problem with the setup

OpenStudy (anonymous):

do you have the answer?

OpenStudy (anonymous):

It's been a while since I did these.

OpenStudy (anonymous):

I have some clues The first is that since I'm rotating around a vertical line, the cross section is going to be a disk

OpenStudy (anonymous):

I think the answer is possibly \[\int\limits_{-1}^{1} (2y^2-2)^2 dy \]

OpenStudy (anonymous):

but no gaurantees

OpenStudy (anonymous):

there should be a pi

OpenStudy (anonymous):

The center of the disk is at x=2, so the radius is 2-2y^2, but I don't under stand how this was arrived at

OpenStudy (anonymous):

I am fairly sure it doesn't matter if it is (2-2y^2) or (2y^2-2) as you sqaure the expression anyway.

OpenStudy (anonymous):



OpenStudy (anonymous):

So I have to figure out A(x) and take the integral of that but I'm lost on how to get the area of the horizontal slice

OpenStudy (anonymous):

Not the best. You need the length of the rectangle ( which is the distance between x=2 and the graph) , how do you find the find the distance between two points, you substract them

OpenStudy (anonymous):

Uh uh

OpenStudy (anonymous):

on the graph, the x coordinate is x=2y^2

OpenStudy (anonymous):

Then that distance represents the radii of the little discs?

OpenStudy (anonymous):

How do you find the area of a disc, pi r^2

OpenStudy (anonymous):

But this region bounded by the function and x=2 is rotated around x=2

OpenStudy (anonymous):

then once you have the area you multiply by "dy" to get the volume.

OpenStudy (anonymous):

so the radius in this case is pi(2y^2)^2?

OpenStudy (anonymous):

no, the radius is (2y^2-2)

OpenStudy (anonymous):

Then the area of the discs is = pi (2y^2-2)^2

OpenStudy (anonymous):

Oh, right, so A(x) = pi(2y^2-2)^2?

OpenStudy (anonymous):

yes, except I think you mean A(y)

OpenStudy (anonymous):

Yes, A(y), sorry

OpenStudy (anonymous):

As it is rotated about the line x=2, you intersect the graph with x=2 and these give you the bounds on y , the limits which you should integrate over.

OpenStudy (anonymous):

-1,1? because that's where 2y^2 intersects with x=2

OpenStudy (anonymous):

\[ Vol = \pi \int\limits_{-1}^{1} [2(y^2-1)]^2 dy\]

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

\[Vol = 4 \pi \int\limits_{-1}^{1} (y^2-1)^2dy\]

OpenStudy (anonymous):

Should I not square the function before taking the integral?

OpenStudy (anonymous):

\[Vol = 8 \pi \int\limits_{0}^{1} (y^2-1)^2 dy\] As integrand is even.

OpenStudy (anonymous):

lol im probably confusing you , all I am doing is simiplifying . I did sqaure it above.

OpenStudy (anonymous):

Now I am really confused because (2y^2-2)^2 = 4y^4-4y^2+4 = 4(y^4-y^2+1)

OpenStudy (anonymous):

no, (2y^2 -2)^2 = 4y^4 -8y^2 +4

OpenStudy (anonymous):

Oh, duh

OpenStudy (anonymous):

That's why I like to simplify it, minimises mistakes. I factored out the 2 and sqaure it by itself

OpenStudy (anonymous):

That's where the 4 comes from at the front of the integral, 2^2 .

OpenStudy (anonymous):

But technically, I could integrate it without simplifying, just dragging pi out front

OpenStudy (anonymous):

And then just simplify at the end when I plug in the limits of integration

OpenStudy (anonymous):

Anyway, \[Vol= 8 \pi \int\limits_{0}^{1} (y^4 -2y^2 +1) = 8 \pi ( \frac{1}{5} -\frac{2}{3}+1) = \frac{64 \pi}{15}\]

OpenStudy (anonymous):

Web assign asks "The cross-sectional area of the disk is A = π( ___). When I plugged in 4y^4-8y^2+4 it gave me an error

OpenStudy (anonymous):

It wants me to do these in pieces